OpenStudy (askme12345):

If lnA=10, write ln(Ae^3) as an integer

OpenStudy (jdoe0001):

keep in mind that \(\bf log(a\times b) \implies log(a) + log(b)\\\quad \\ ln = log_e\)

OpenStudy (jdoe0001):

also keep in mind the logarithm cancellation rule of \(\bf \large log_aa^x = x\)

OpenStudy (askme12345):

so is the correct answer 10^3 ... I have these rules but i am not following thank you

OpenStudy (jdoe0001):

ahemm... no

OpenStudy (jdoe0001):

try using the \(\bf log(a\times b) \implies log(a) + log(b)\) see what does that give you

OpenStudy (jdoe0001):

keep in mind also that \(\bf ln(Ae^3) \implies ln(A \times e^3)\)

OpenStudy (askme12345):

would it be possible for you to assist me with solving the answer? I see the rules, but am having a hard time applying them to the question

OpenStudy (jdoe0001):

hmmm.... lemme see if this helps \(\bf log_e(a\times b) \implies log_e(a) + log_e(b)\\\quad \\ ln(Ae^3) \implies log_e(Ae^3) \implies log_e(A \times e^3)\)

OpenStudy (askme12345):

can u plug in the numbers for me, and give me a visual please i am still clueless